1. Which ordered pair is the solution of the following system of equations?
3x+2y=4
-2x+2y=24

Answer by richwmiller(17219)

3x+2y=4
-2x+2y=24
subtract the second from the first
5x=-20
x=-4
3*(-4)+2y=4
-12+2y=4
2y=16
y=8
(-4,8)
check on second equation
-2(-4)+(2*8)=24
8+16=24
24=24 ok
check first equation
3*(-4)+2*8=4
-12+16=4
4=4 ok